# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v2

@@ 4.1. Resistance of the series-parallel graph
$g(s)=\left\{\begin{array}[]{r l}\log(e^{s}-1),&\text{if}\,\,s\geq\log 2,\\ 0,&\text{otherwise.}\end{array}\right$.
@@
In this case, the constant $a$ is given by $a=2\zeta(3)$. See Appendix C for the proof.
@@
$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}),$
where $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
@@
Let $X^{(n)}=\log R^{(n)}$. As in the discussion in the introduction, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$. Let $q_{-}=\mathbb{P}\{X^{(0)}=-\infty\}$ and $q_{+}=\mathbb{P}\{X^{(0)}=+\infty\}$ so that $1-q=q_{+}+q_{-}$. Notice that
@@
Let $F_{N}:\mathbb{R}\times[0,+\infty)\to[0,1]$ be the rescaled CDF of $\{X^{(n)}\}$ defined as in Theorem 2. By that theorem, $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
@@
$\partial_{t}G-q^{2}a|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
@@
$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qa)^{\frac{1}{3}}}+\frac{1}{2}.$
@@
$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,t)-q_{-})=G(x,t).$
@@ 4.2. Distance on the series-parallel graph
A complete analysis of the asymptotics of this RDE in the subcritical case ($p<\frac{1}{2}$) is given in [21]. There is a random variable $K$ such that if $\lambda$ denotes the essential infimum of $D^{(0)}$, then $D^{(n)}\to\lambda K$ as $n\to\infty$. This implies that $K$ is the unique (up to multiplication by a constant) fixed point of the RDE in this case.
In the critical case ($p=\frac{1}{2}$), the main result of [6] obtains a $\text{Beta}(2,1)$ limit for $N^{-1/2}\log D^{(N)}$ provided $D^{(0)}=1$ almost surely. The next corollary of Theorem 1 describes the asymptotics for an arbitrary initial distribution. Notice the contrast with the subcritical case: In the subcritical case, if $D^{(0)}$ can be arbitrarily small with positive probability, then $D^{(n)}\to 0$ as $n\to\infty$, whereas here, conditionally on the event $\{D^{(n)}>0\}$, mass at zero only slows the rate of growth.
@@
Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<+\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$. Then $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
@@
The claimed value of $\sigma$ is computed in Appendix C and is consistent with [6]. The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$. ∎
@@ 4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
The RDE associated with this equation provides a very natural distribution-dependent random walk on $\mathbb{Z}$, and it has the nice feature that the equations describing the CDF are nothing other than finite-difference discretizations of the quasilinear Fisher-KPP equations mentioned in the introduction.
@@
where $g_{\mathbb{Z}}$ denotes the left-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$. Thus, such a law $\mathbf{P}$ satisfies the assumptions of Theorems 1 and 2. In particular, as long as $(f_{+},f_{-})$ are not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
@@
(38) $F_{n+1}-F_{n}$ $=-pq_{+}(\nabla_{\mathbb{Z}}^{-}F_{n})^{2}+(1-p)q_{-}(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}),$
where $\nabla_{\mathbb{Z}}^{+}$, $\nabla_{\mathbb{Z}}^{-}$ are the forward- and backward-difference operators
@@
(39) $F_{n+1}-F_{n}$ $=a\nabla^{\text{sym}}_{\mathbb{Z}}F_{n}\Delta_{\mathbb{Z}}F_{n}+\sigma(\nabla_{\mathbb{Z}}^{+}F_{n})^{2}+(1-2p)F_{n}(1-F_{n}).$
@@
Equation (39) looks like a discrete analogue of the quasilinear Fisher-KPP equation (24) derived above. Notice that here $a$ or $\sigma$ may vanish, irrespective of whether or not $p=\frac{1}{2}$. Since replacing $\mathbb{Z}$ by $\mathbb{R}$ and finite-differences by derivatives does not seem likely to fundamentally change the large-scale asymptotic behavior, equation (39) motivates the general study of (24) for arbitrary $\sigma\in\mathbb{R}$ and $a\geq 0$.
It is worth emphasizing that there is no contradiction between the formulas (39) and (33) since the former involves the CDF of a $\mathbb{Z}$-valued random variable while the latter describes smooth CDF’s.
@@ 4.4. Hipster Random Walks and Cooperative Motion
Inspection of (39) also reveals a connection with the symmetric hipster random walk and two-player cooperative motion, examples considered in in [1] and [3], respectively.
To understand the connection, consider again the random variable $Z$ defined by $Z=\Theta\max\{X_{1},X_{2}\}+(1-\Theta)\min\{X_{1},X_{2}\}$. Recall that $Z$ and $X_{1}$ have the same distribution if $p=\frac{1}{2}$. At the same time, since the two coincide when $X_{1}=X_{2}$, a stronger statement is true: $Z$ and $X_{1}$ have the same conditional laws on the event $\{X_{1}\neq X_{2}\}$. The same reasoning applies if $X_{1}$ is replaced by the more symmetric $\Theta X_{1}+(1-\Theta)X_{2}$. From this, it follows that the law of the output $Y$ is unchanged if the equation is instead changed to
@@
Note that the equation $q=\frac{1}{2}q_{+}$ introduces the constraint $q\leq\frac{1}{2}$, so that the previous result only covers the asymmetric hipster random walk for half of the possible values of $q$. In fact, it is possible to show (by manipulation of the evolution equation satisfied by the CDF) that the asymmetric hipster random walk is a monotone RDE only if $q\leq\frac{1}{2}$. Therefore, it cannot be described using (9) outside this regime.
@@
The RDE is parameterized by the pair $(r,q)$. The next result describes the connection with the RDE’s considered here. As for the asymmetric hipster random walk, only a subset of the parameter values are covered.
@@
(46) $\{(r,q)\,\mid\,q\in[0,1],\,\,|r-\frac{1}{2}|\leq\frac{1}{2}(q^{-1}-1)\}.$
@@
In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main results. The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above. Interestingly, in the asymmetric case, those works were able to obtain limit theorems even in the nonmonotone regime.
@@ 5. Scaling Limit
Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define a sequence $\{p^{(N)}\}$ by
$p^{(N)}=1/2+\theta N^{-1}.$
Let $T^{(N)}$ be the operator associated with the RDE (9) with $\Phi$ given by (8) and bias parameter $p=p^{(N)}$. In view of the results of Section 3, the equation $F_{n}=T^{(N)}F_{n-1}$ can be written in the form
(47) $F_{n}-F_{n-1}=\mathcal{L}^{(N)}F_{n-1}-2\theta N^{-1}F_{n-1}(1-F_{n-1})$
@@ 5.1. Abstract Framework
    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
@@
    Consistency of $\mathcal{L}^{(N)}$: If $\Phi\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $\Phi_{\delta}(y)=\Phi(\delta y)$, then
    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}\Phi_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}\Phi(x),\partial_{x}^{2}\Phi(x)),$
@@
    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to[0,1]$ such that
@@
$F^{(N)}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
@@
$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,\infty)\,\,\text{as}\,\,N\to+\infty,$
where $F$ is the unique bounded viscosity solution of the equation
(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
It is possible to show that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details. In particular, $\{Q^{(N)}\}$ is given by $Q^{(N)}(q)=2\theta N^{-1}q(1-q)$, which trivially satisfies (vi). It remains to verify condition (v). That is the subject of the next proposition. Recall that the constants $\sigma$ and $a$ are defined in Section 1.4.1 above.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and fix $\theta\in\mathbb{R}$. Define $p^{(N)}=1/2+\theta N^{-1}$. Consider the operator $\mathscr{L}^{(N)}$ given by (32) with $p=p^{(N)}$. If $\Phi:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $\Phi_{\delta}(y)=\Phi(\delta y)$, and if $R_{\Phi}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined by
$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}R_{\Phi}^{(N)}(x),$
then $R^{(N)}_{\Phi}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
Further, if the stronger assumption (16) holds, and if $R^{(N)}_{\Phi}$ is instead defined by
$(\mathscr{L}^{(N)}\Phi_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}\Phi(x)|^{2}+N^{-1}a|\partial_{x}\Phi(x)|\partial_{x}^{2}\Phi(x)+N^{-1}R_{\Phi}^{(N)}(x),$
then, as before, $R^{(N)}_{\Phi}\to 0$ uniformly as $N\to+\infty$.
@@
    Asymptotically regularizing: If $\overline{F}$ and $\underline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
    $\overline{F}=\underline{F}\quad\text{in}\,\,\mathbb{R}\times(0,\infty).$
As explained in Appendix A, from the fact that $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
@@
It follows from classical results that, in the PDE’s of interest here, assumption (vii) holds as long as either $\sigma\neq 0$ or $a>0$. (The degenerate PDE’s $\partial_{t}F-Q(F)=0$ do not have this property and therefore require a slightly different treatment.) This is made precise in the next proposition.
@@
Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\geq 0$ and $Q(1)\leq 0$. If $\mathcal{F}(v,w)=\sigma|v|^{2}$ for some $\sigma\neq 0$ or $\mathcal{F}(v,w)=a|v|w$ for some $a>0$, then assumption (vii) is satisfied.
@@
Theorem 3 is stated for semigroups defined in $CDF(\overline{\mathbb{R}})$, the space of nondecreasing, right-continuous functions taking values in $[0,1]$. There is nothing special about right-continuity, however. In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-increasing versions of a nondecreasing function $F$. If operators $\{T^{(N)}_{-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T^{(N)}_{-}F=(T^{(N)}(F^{+}))_{-}$, then $\{T^{(N)}_{-}\}$ satisfies assumptions (i)-(vi) if and only if $\{T^{(N)}\}$ does. Further, it is not hard to see that a PDE scaling limit of $\{T^{(N)}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T^{(N)}_{-}\}$.

### v3

@@ 4.1. Resistance of the series-parallel graph
$g(s)=\log(e^{s}-1)\quad\text{for each}\,\,s\geq\log 2.$
@@
By symmery, $\sigma=0$ in this setting and $a$ can be explicitly computed to be $a=2\zeta(3)$. See Appendix C for the proof.
@@
$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}).$
Here $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
@@
Let $X^{(n)}=\log R^{(n)}$. As in the discussion above, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$. Let $q_{-}=\mathbb{P}\{X^{(0)}=-\infty\}$ and $q_{+}=\mathbb{P}\{X^{(0)}=+\infty\}$ so that $1-q=q_{+}+q_{-}$. Notice that
@@
Let $F_{N}:\mathbb{R}\times[0,+\infty)\to[0,1]$ be the rescaled CDF of $\{X^{(n)}\}$ defined as in Theorem 2. By that theorem (which applies even though $\{X^{(n)}\}$ are extended real-valued, see Remark 3 below), $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
@@
$\partial_{t}G-qa|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
@@
$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qat)^{\frac{1}{3}}}+\frac{1}{2}.$
@@
$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,1)-q_{-})=G(x,1).$
@@ 4.2. Distance on the series-parallel graph
A complete analysis of the asymptotics of this RDE in the subcritical case ($p<1/2$) was given in [21]: There is a random variable $K$ such that if $\lambda$ denotes the essential infimum of $D^{(0)}$, then $D^{(n)}\to\lambda K$ as $n\to\infty$. This implies that $K$ is the unique (up to multiplication by a constant) fixed point of the RDE in this case.
At the critical point $p=1/2$, the main result of [6] obtains a $\text{Beta}(2,1)$ limit for $N^{-1/2}\log D^{(N)}$ provided $D^{(0)}=1$ almost surely. The next corollary of Theorem 1 describes the asymptotics for an arbitrary initial distribution, allowing for the possibility that the i.i.d. edge weights can be zero or $\infty$. Notice the contrast with the subcritical case: In the subcritical case, if $D^{(0)}$ can be arbitrarily small with positive probability, then $D^{(n)}\to 0$ as $n\to+\infty$, whereas here, conditionally on the event $\{0<D^{(n)}<+\infty\}$, mass at zero only slows the rate of growth.
@@
Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}\leq+\infty\}=1$ and let $q=\mathbb{P}\{0<D^{(0)}<+\infty\}$. Then $\mathbb{P}\{0<D^{(n)}<+\infty\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,0<D^{(N)}<+\infty\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
@@
The claimed value of $\sigma$ is computed in Appendix C and is consistent with [6]. The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{0<D^{(n)}<+\infty\}=q$ for any $n$ by Proposition 7. ∎
@@ 4.3. $\mathbb{Z}$-valued Solutions and Finite-Difference Equations
The RDE associated with this equation provides a very natural distribution-dependent random walk on $\mathbb{Z}$, and it has the nice feature that the equations describing the CDF are nothing other than finite-difference discretizations of the quasilinear Fisher-KPP equations discussed in the introduction above.
@@
where, as in Section 1.4.1, $g_{\mathbb{Z}}$ denotes the right-continuous, right-inverse of the function $u\mapsto u+f_{\mathbb{Z}}(u)$. Thus, such a law $\mathbf{P}$ satisfies the assumptions of Theorems 1 and 2. In particular, as long as $(f_{+},f_{-})$ are not identically zero, either $n^{-\frac{1}{2}}X^{(n)}$ or $n^{-\frac{1}{3}}X^{(n)}$ converges to a nondegenerate limit.
@@
(38) $F_{n}-F_{n-1}$ $=-pq_{+}(\nabla_{\mathbb{Z}}^{-}F_{n-1})^{2}+(1-p)q_{-}(\nabla_{\mathbb{Z}}^{+}F_{n-1})^{2}+(1-2p)F_{n-1}(1-F_{n-1}),$
where $\nabla_{\mathbb{Z}}^{+}$ and $\nabla_{\mathbb{Z}}^{-}$ are the forward- and backward-difference operators
@@
(39) $F_{n}-F_{n-1}$ $=a\nabla^{\text{sym}}_{\mathbb{Z}}F_{n-1}\Delta_{\mathbb{Z}}F_{n-1}+\sigma(\nabla_{\mathbb{Z}}^{+}F_{n-1})^{2}+(1-2p)F_{n-1}(1-F_{n-1}).$
@@
Equation (39) looks like a discrete analogue of the quasilinear Fisher-KPP equation (24) discussed in Section 1.5. Notice that here $a$ or $\sigma$ may vanish, irrespective of whether or not $p=1/2$. Since replacing $\mathbb{Z}$ by $\mathbb{R}$ and finite-differences by derivatives does not seem likely to fundamentally change the large-scale asymptotic behavior, equation (39) motivates the general study of (24) for arbitrary $\sigma\in\mathbb{R}$ and $a\geq 0$.
It is worth emphasizing that there is no contradiction between the formulae (39) and (33) since the former involves the CDF of a $\mathbb{Z}$-valued random variable while the latter describes smooth CDF’s.
@@ 4.4. Hipster Random Walks and Cooperative Motion
Inspection of (39) also reveals a connection with the symmetric hipster random walk and two-player cooperative motion, examples considered in [1] and [3], respectively. For the rest of this section, let $p=1/2$.
To understand the connection, consider again the random variable $Z$ defined by $Z=\Theta\max\{X_{1},X_{2}\}+(1-\Theta)\min\{X_{1},X_{2}\}$. Recall that $Z$ and $X_{1}$ have the same distribution since $p=1/2$. At the same time, in view of the fact that the two coincide when $X_{1}=X_{2}$, a stronger statement is true: $Z$ and $X_{1}$ have the same conditional laws on the event $\{X_{1}\neq X_{2}\}$. This reasoning also applies if $X_{1}$ is replaced by the more symmetric $\Theta X_{1}+(1-\Theta)X_{2}$. From this, it follows that the law of the output $Y$ from (37) is unchanged if the equation is instead changed to
@@
Note that the equation $q=\frac{1}{2}q_{+}$ introduces the constraint $q\leq\frac{1}{2}$, so that the previous result only covers the asymmetric hipster random walk for half of the possible values of $q$. In fact, it is possible to show (by manipulation of the evolution equation satisfied by the CDF) that the asymmetric hipster random walk is a monotone RDE only if $q\leq 1/2$. Therefore, it falls outside the class of RDE’s considered in this paper when $q>1/2$.
@@
The RDE is parameterized by the pair $(r,q)$. The next result describes the connection with the RDE’s considered here. As for the asymmetric hipster random walk, only a subset of the parameter values are covered; the proof is left to the interested reader.
@@
(46) $\left\{(r,q)\,\mid\,q\in[0,1],\,\,\left|r-\frac{1}{2}\right|\leq\frac{1}{2}(q^{-1}-1)\right\}.$
@@
In view of the proposition above, the $\text{Beta}(2,2)$ limit theorems for the symmetric RDE’s considered in [1] and [3] follow as a special case of this paper’s main results. The $\text{Beta}(2,1)$ limit theorems for asymmetric RDE’s obtained in [1] and [2] also follow, albeit only in the parameter regimes identified above. Interestingly, in the asymmetric case, those works were able to prove limit theorems even in the nonmonotone regime.
@@ 5. Scaling Limit
Fix data $(f_{+},f_{-})$ satisfying the assumptions of Theorems 1 or 2. Given $\theta\in\mathbb{R}$, define the sequence $\{p^{(N)}\}$ by
$p^{(N)}=\frac{1}{2}+\theta N^{-1}.$
Let $T^{(N)}$ be the operator associated with the RDE (9) with $\Phi$ given by (8) and bias parameter $p=p^{(N)}$. In view of the results of Section 3, the equation $F^{(N)}_{n}=T^{(N)}F^{(N)}_{n-1}$ can be written in the form
(47) $F^{(N)}_{n}-F^{(N)}_{n-1}=\mathcal{L}^{(N)}F^{(N)}_{n-1}-2\theta N^{-1}F^{(N)}_{n-1}(1-F^{(N)}_{n-1})$
@@ 5.1. Abstract Framework
    Discrete-Time Reaction-Diffusion Equation: There are functions $\{\mathcal{L}^{(N)}\}$ defined on $CDF(\overline{\mathbb{R}})$ and continuous functions $\{Q^{(N)}\}$ defined on $[0,1]$ such that
@@
    Consistency of $\mathcal{L}^{(N)}$: If $G\in CDF(\overline{\mathbb{R}})$ is smooth with bounded, uniformly continuous second derivatives, rescaled according to the formula $G_{\delta}(y)=G(\delta y)$, then
    $\lim_{N\to+\infty}N(\mathscr{L}^{(N)}G_{\delta^{(N)}})\left(\frac{x}{\delta^{(N)}}\right)=\mathcal{F}(\partial_{x}G(x),\partial_{x}^{2}G(x)),$
@@
    Consistency of $Q^{(N)}$: There is a function $Q:[0,1]\to\mathbb{R}$ such that
@@
$F_{N}(\cdot,0)\to F_{\text{in}}\quad\text{vaguely as}\,\,N\to+\infty,$
@@
$F_{N}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times[0,+\infty)\,\,\text{as}\,\,N\to+\infty,$
where $F$ is the unique bounded (continuous) viscosity solution of the equation
(49) $\left\{\begin{array}[]{r l}\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)=0&\text{in}\,\,\mathbb{R}\times(0,+\infty),\\ F(x,0)=F_{\text{in}}(x).\end{array}\right$.
The results of the previous sections imply that if $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, as is assumed in Theorems 1 and 2, then the one-parameter family $\{T^{(N)}\}$ defined in the discussion at the beginning of this section satisfies assumptions (i)-(iv); see Section 5.2 for the details. In particular, $\{Q^{(N)}\}$ is given by $Q^{(N)}(q)=2\theta N^{-1}q(1-q)$, which trivially satisfies (vi). It remains to verify condition (v). That is the subject of the next proposition. Recall that the constants $\sigma$ and $a$ are defined in Section 1.4.1 above.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=1$ and (15) holds, and fix $\theta\in\mathbb{R}$. Define $p^{(N)}=1/2+\theta N^{-1}$. Consider the operator $\mathscr{L}^{(N)}$ given by (32) with $p=p^{(N)}$. If $G:\mathbb{R}\to[0,1]$ is a nondecreasing function with bounded, uniformly continuous first and second derivatives, rescaled according to the rule $G_{\delta}(y)=G(\delta y)$, and if $R_{G}^{(N)}:\mathbb{R}\to\mathbb{R}$ is the remainder defined so that
$(\mathscr{L}^{(N)}G_{N^{-\frac{1}{2}}})(N^{\frac{1}{2}}x)=N^{-1}\sigma|\partial_{x}G(x)|^{2}+N^{-1}R_{G}^{(N)}(x),$
then $R^{(N)}_{G}\to 0$ uniformly in $\mathbb{R}$ as $N\to+\infty$.
Further, if the stronger assumption (16) holds, and if $R^{(N)}_{G}$ is instead defined by
$(\mathscr{L}^{(N)}G_{N^{-\frac{1}{3}}})(N^{\frac{1}{3}}x)=N^{-\frac{2}{3}}\sigma|\partial_{x}G(x)|^{2}+N^{-1}a|\partial_{x}G(x)|\partial_{x}^{2}G(x)+N^{-1}R_{G}^{(N)}(x),$
then, as before, $R^{(N)}_{G}\to 0$ uniformly as $N\to+\infty$.
@@
    Asymptotically regularizing: If $\underline{F}$ and $\overline{F}$ denote the maximal and minimal viscosity sub- and supersolutions, respectively, of the initial-value problem (49) for a given $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$ (see Appendix A.1), then
    $\underline{F}=\overline{F}\quad\text{in}\,\,\mathbb{R}\times(0,+\infty).$
As explained in Appendix A, from the fact that $\underline{F}=\overline{F}$ in $\mathbb{R}\times(0,+\infty)$, it follows that there is a unique solution of (49) started from any initial datum $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$, which is continuous for positive times.
@@
It follows from classical results that, in the PDE’s of interest here, assumption (vii) holds as long as either $\sigma\neq 0$ or $a>0$. (The degenerate PDE $\partial_{t}F+Q(F)=0$ does not have this property and therefore requires a slightly different treatment.) This is made precise in the next proposition.
@@
Let $Q$ be an arbitrary function for which there is an $L>0$ such that $|Q(q)-Q(q^{\prime})|\leq L|q-q^{\prime}|$ for each $q,q^{\prime}\in[0,1]$ and satisfying both $Q(0)\leq 0$ and $Q(1)\geq 0$. If $\mathcal{F}(v,w)=\sigma|v|^{2}$ for some $\sigma\neq 0$ or $\mathcal{F}(v,w)=a|v|w$ for some $a>0$, then assumption (vii) is satisfied.
@@
Theorem 3 is stated for semigroups defined in $CDF(\overline{\mathbb{R}})$, the space of nondecreasing, right-continuous functions taking values in $[0,1]$. There is nothing special about right-continuity, however. In fact, let $F_{-}$ and $F^{+}$ denote the left- and right-continuous versions of a nondecreasing function $F$. If operators $\{T^{(N)}_{-}\}$ are defined on the space $CDF_{\text{left}}(\overline{\mathbb{R}})$ of nondecreasing, left-continuous functions taking values in $[0,1]$ by the formula $T^{(N)}_{-}F=(T^{(N)}(F^{+}))_{-}$, then $\{T^{(N)}_{-}\}$ satisfies assumptions (i)-(vi) if and only if $\{T^{(N)}\}$ does. Further, it is not hard to see that a PDE scaling limit of $\{T^{(N)}\}$ holds as in Theorem 3 if and only if the same conclusion applies to $\{T^{(N)}_{-}\}$.